Another physics problem. Please help.
James Joule (after whom the unit of energy is name
Answer #1

Potential Energy of 1 kg of water at 51m from ground:

U = Mgh = 1(9.8)(51) = 499.8 J

At the bottom, this will be converted to thermal energy.

499.8 = mc\Delta T

=> 499.8 = (1)(4190)\Delta T

=> \Delta T = \frac{499.8}{(1)(4190)} = 0.12 K = 0.12 \degree C